OpenStudy (anonymous):

Which rule will find the nth term of the arithmetic sequence -58, -65, -72, -86, …?

OpenStudy (anonymous):

looks like you are subtracting \(7\) each time

OpenStudy (dan815):

Hi satellite :)

OpenStudy (anonymous):

try \(-58-7(n-1)\) could work there are other ways to write it

OpenStudy (anonymous):

an = -65 − 7n an = -65 + 7n an = 51 − 7n an = -51 + 7n an = -51 − 7n

OpenStudy (anonymous):

yall i am on a test i have 7 tests to take in 2 hours pleeeeeeeeeeeeeeeeeeez

OpenStudy (anonymous):

multiply out \[-58-7(n-1)\]\[-57-7n+7\] \[-51-7n\]

OpenStudy (anonymous):

thank you.